Session: Feb 2000 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Mon Feb 14 14:00:08 MET 2000) Success
No known problems.

Medicina (Franco Mantovani at Mon Feb 14 14:11:20 MET 2000) Success
No problems during the observation. Clear sky.

Effelsberg (A. Kraus at Tue Feb 15 09:17:09 MET 2000) Minor failures
All scans between UT0700-0830 had been lost due to tape problems.

Jodrell (Alastair Gunn at Wed Feb 16 05:48:19 MET 2000) Success
PI did not allow adequate slewing times for some scans, particularly at start of experiment. No other known problems.

Noto (C.Trigilio at Thu Feb 17 12:51:41 MET 2000) Success

Westerbork (Tony Foley at Thu Feb 17 17:35:05 MET 2000) Minor failures
Stopped early (12:03UT) as all scans after that were out of Hour Angle range. No other known problems

Calibration (Alastair Gunn at Wed Apr 19 13:00:56 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/gk020.antab for details